# Satellite Federated Fine-Tuning for Foundation Models in Space Computing Power Networks

### v1

@@ II-A Centralized Satellite FL
In centralized satellite FL, a coordinator, such as a PS, collects model updates from satellites. Satellites perform local training on their datasets and then send the model updates to the central coordinator. Some algorithms employ only SGLs during training, where models are trained on-board and transmitted via SGLs for global model aggregation at the PS. These algorithms are usually designed to address the challenges of time-varying and unstable SGLs. For instance, the algorithm proposed in [15] adjusted aggregation intervals based on the density of satellite-ground connections to increase the frequency of model updates.  [16] proposed an over-the-air computation based scheme and optimized both uplink and downlink communications through beamforming. Considering the variations in connection time and connection density among satellites, asynchronous FL is introduced to mitigate the effects of stragglers, allowing satellites to send model updates to the server asynchronously [17]. A compensation mechanism proposed in [18] addressed gradient staleness by adjusting the updates according to discrepancies.
@@ III-C Satellite Federated Fine-Tuning for Remote Sensing Foundation Models
Remote sensing foundation models (RS FMs) are specifically designed to analyze RS data from remote sensing devices, such as satellites and drones. These huge FMs are pre-trained and can be fine-tuned to enhance their accuracy and efficiency in specific downstream tasks. Typically, RS FMs consist of the embedding layer, backbone network, and task-specific head, as shown in Fig. 1. Take SpectralGPT as an example, the workflow is as follows [9]. Images are split into patches and the embedding layer embeds image patches, adds position embeddings, and outputs embedding vectors. Then, the embedding vectors are fed into the backbone network to extract feature vectors. The computation load of the backbone network accounts for the vast majority of the total computation load of the model. Finally, for each downstream task, a learnable task-specific head is connected to the model. This head receives feature vectors as inputs and outputs results. During fine-tuning, only head parameters ${\bm{w}}^{H}$ are updated, while other parameters remain fixed in each training round $r$.
@@
where $\eta$ represents the learning rate, and $\nabla\operatorname{L}()$ denotes the gradient of the loss function. Since only the head parameters are updated, only these parameters need to be aggregated:
@@
Our primary objective is to address the computing power constraints of fine-tuning FMs on satellites. To achieve this, we employ a computation task decomposition technique within our framework, effectively splitting the FM into distinct components and deploying them based on their computation requirements and distributed computation network. To protect data privacy, we deploy the embedding layer and the task-specific head on the satellites. The backbone network, which bears most of the computational workload (approximately 99%), is deployed on the ground. This significantly reduces the computational burden on the satellites, allowing them to handle only about $1\text{\textperthousand}$ of the overall computational tasks. During the forward propagation process, embedding vectors and feature vectors are exchanged between the satellites and the ground. Notably, the size of these vectors is approximately 20 MB for a sample, a reduction of about a thousand times compared to the raw satellite imagery. This drastic reduction in data size not only enhances communication efficiency but also alleviates bandwidth requirements, making the system more cost-effective and practical for deployment in space-constrained environments.
@@ IV-A1 Intra-Orbit Embedding Vector Broadcast
First, each satellite $S_{p,n}$ computes embedding vectors $\bm{E}_{p,n}$ utilizing data stored locally. Then, all computed embedding vectors are broadcast within each orbit so that any satellite connected to a GS can transmit intra-orbit data to the ground for forward propagation. This transmission scheme significantly reduces waiting times compared to the conventional transmission approach where each satellite only sends its local embedding vectors to the ground. We visualize the connection windows between satellites and GSs throughout the day in Fig. 5 for comparison. Each connection window lasts for approximately $5\sim 10$ minutes. If data transmission starts at 23:30, some satellites establish their first satellite-ground communication around three hours later. So conventional methods waste the communication resources of other satellites. In contrast, our proposed method allows other intra-orbit satellites to transmit data immediately without waiting.
@@ IV-B2 Intra-orbit Head Aggregation
After local fine-tuning, the updated task-specific heads from all satellites within the same orbit are averaged to ensure model convergence in collaborative learning. Satellites exchange information with each other in a decentralized way to aggregate these local heads. This head aggregation process does not rely on a terrestrial coordinator and accounts for the intermittent nature of satellite-ground communication. To obtain the global model, we perform hierarchical aggregation through ISLs. Head parameters are first aggregated within each orbit
@@
and then aggregated across different orbits. Due to the more stable communication conditions maintained by satellites within the same orbit, intra-orbit head aggregation employs a parallel communication strategy similar to that used in the embedding vector broadcast. This approach ensures efficient and lower-cost transmission times.
@@ VII Conclusion
This paper proposed a satellite federated fine-tuning framework for RS FMs. The proposed framework addresses the training latency and computation resource constraints stemming from the limited computational capabilities of satellites, intermittent SGLs, and dynamic ISLs in Space-CPN. To tackle these issues, we designed a model partitioning scheme and training process. We partition models into the embedding layer, the backbone network and the output layer and deploy the backbone networks and other components on the PS and satellites respectively to reduce on-board computation load and transmission burden. Furthermore, communication strategies for both satellite-ground and inter-satellite transmissions were co-designed with the training process to minimize latency. We put forward a topology-aware communication scheme, which assigns transmission tasks according to the real-time topological structure and the status of links. Regarding intra-orbit communication, considering the ring topology of each orbit, we present a parallel communication method founded on Ring Allreduce. As for inter-orbit communication, we design a communication algorithm that aims to reduce latency to the minimum by taking link capacity into account. These strategies blend communication and computational procedures, factoring in aspects such as restricted bandwidth, sparse connectivity, and dynamic topologies. As a result, these strategies enhance the efficiency of data transmission and expedite the convergence of models. Simulation results validate the effectiveness of our proposed framework and communication strategies.

### v2

@@ II-A Centralized Satellite FL
In centralized satellite FL, a coordinator, such as a PS, collects model updates from satellites. Satellites perform local training on their datasets and then send the model updates to the central coordinator. Some algorithms employ only SGLs during training, where models are trained on-board and transmitted via SGLs for global model aggregation at the PS. These algorithms are usually designed to address the challenges of time-varying and unstable SGLs. For instance, the algorithm proposed in [15] adjusted aggregation intervals based on the density of satellite-ground connections to increase the frequency of model updates. [16] proposed an over-the-air computation based scheme and optimized both uplink and downlink communications through beamforming. Considering the variations in connection time and connection density among satellites, asynchronous FL is introduced to mitigate the effects of stragglers, allowing satellites to send model updates to the server asynchronously [17, 18]. A compensation mechanism proposed in [19] addressed gradient staleness by adjusting the updates according to discrepancies.
@@ III-C Satellite Federated Fine-Tuning for Remote Sensing Foundation Models
Remote sensing foundation models (RS FMs) are specifically designed to analyze RS data from remote sensing devices, such as satellites and drones. These huge FMs are pre-trained and can be fine-tuned to enhance their accuracy and efficiency in specific downstream tasks. Typically, RS FMs consist of the embedding layer, backbone network, and task-specific head, as shown in Fig. 1. Take SpectralGPT as an example, the workflow is as follows [9]. First, images are split into patches and the embedding layer embeds image patches, adds position embeddings, and outputs embedding vectors. Then, the embedding vectors are fed into the backbone network to extract feature vectors. The computation load of the backbone network accounts for the vast majority of the total computation load of the model. Finally, for each downstream task, a learnable task-specific head is connected to the model. This head receives feature vectors as inputs and outputs results. During fine-tuning, only head parameters ${\bm{w}}^{H}$ are updated, while other parameters remain fixed in each training round $r$.
@@
where $\eta$ represents the learning rate, and $\nabla\operatorname{L}()$ denotes the gradient of the loss function. Since only the head parameters are updated, only head parameters need to be aggregated:
@@
Our primary objective is to address the computing power constraints of fine-tuning FMs on satellites. To achieve this, we employ a computation task decomposition technique, effectively splitting the FM into distinct components and deploying them based on requirements. To protect data privacy, we deploy the embedding layer and the task-specific head on the satellites. The backbone network, which bears most of the computational workload (approximately 99%), is deployed on the ground. This significantly reduces the computational burden on the satellites, allowing them to handle only about $1\text{\textperthousand}$ of the overall computational tasks. During the forward propagation process, embedding vectors and feature vectors are exchanged between the satellites and the ground. Notably, the size of these vectors is approximately 20 MB for a sample, a reduction of about a thousand times compared to the raw satellite imagery. This drastic reduction in data size not only enhances communication efficiency but also alleviates bandwidth requirements, making the system more cost-effective and practical for deployment in space-constrained environments.
@@ IV-A1 Intra-Orbit Embedding Vector Broadcast
First, each satellite $S_{p,n}$ computes embedding vectors $\bm{E}_{p,n}$ utilizing data stored locally. Then, all computed embedding vectors are broadcast within each orbit so that any satellite connected to a GS can transmit intra-orbit data to the ground for forward propagation. We visualize the connection windows between satellites and GSs throughout the day in Fig. 5 for comparison. Each connection window lasts for approximately $5\sim 10$ minutes. If data transmission starts at 23:30, some satellites establish their first satellite-ground communication around three hours later. So the conventional method, where each satellite only sends its local embedding vectors without coordination, wastes the communication resources of other satellites. In contrast, our proposed method allows other intra-orbit satellites to transmit data immediately without waiting.
@@ IV-B2 Intra-orbit Head Aggregation
After local fine-tuning, the updated task-specific heads from all satellites within the same orbit are averaged to ensure model convergence in collaborative learning. Satellites exchange head parameters with each other in a decentralized way without a central terrestrial coordinator. Accounting for the intermittent nature of satellite-ground communication, we perform hierarchical aggregation through ISLs. Head parameters are first aggregated within each orbit
@@
and then aggregated across different orbits. Due to the more stable communication conditions maintained by satellites within the same orbit, intra-orbit head aggregation employs a parallel communication strategy similar to that used in the embedding vector broadcast. This approach ensures efficient transmission.
@@ VII Conclusion
This paper proposed a satellite federated fine-tuning framework for RS FMs to addresses the training latency and computation resource constraints stemming from the limited computational capabilities of satellites, intermittent SGLs, and dynamic ISLs in Space-CPN. To tackle these issues, we designed a model partitioning and training scheme. We partition models into the embedding layer, the backbone network and the output layer and deploy the backbone networks and other components on the PS and satellites respectively to reduce on-board computation load and transmission burden. Furthermore, communication strategies for both satellite-ground and inter-satellite transmissions were co-designed with the training process to minimize latency. We put forward a topology-aware communication scheme, which assigns transmission tasks according to the real-time topological structure and the status of links. Regarding intra-orbit communication, considering the ring topology of each orbit, we present a parallel communication method founded on Ring Allreduce. As for inter-orbit communication, we design a communication algorithm that aims to reduce latency to the minimum by taking link capacity into account. These strategies blend communication and computational procedures, factoring in aspects such as restricted bandwidth, sparse connectivity, and dynamic topologies. As a result, these strategies enhance the efficiency of data transmission and expedite the convergence of models. Simulation results validate the effectiveness of our proposed framework and communication strategies.